# Deep Metric Imitation Learning for Stable Motion Primitives

### v2

@@ Abstract
Abstract— Imitation Learning (IL) is a powerful technique for intuitive robotic programming. However, ensuring the reliability of learned behaviors remains a challenge. In the context of reaching motions, a robot should consistently reach its goal, regardless of its initial conditions. To meet this requirement, IL methods often employ specialized function approximators that guarantee this property by construction. Although effective, these approaches come with a set of limitations: 1) they are unable to fully exploit the capabilities of modern Deep Neural Network (DNN) architectures, 2) some are restricted in the family of motions they can model, resulting in suboptimal IL capabilities, and 3) they require explicit extensions to account for the geometry of motions that consider orientations. To address these challenges, we introduce a novel stability loss function, drawing inspiration from the triplet loss used in the deep metric learning literature. This loss does not constrain the DNN’s architecture and enables learning policies that yield accurate results. Furthermore, it is not restricted to a specific state space geometry; therefore, it can easily incorporate the geometry of the robot’s state space. We provide a proof of the stability properties induced by this loss and empirically validate our method in various settings. These settings include Euclidean and non-Euclidean state spaces, as well as first-order and second-order motions, both in simulation and with real robots. More details about the experimental results can be found in: https://youtu.be/ZWKLGntCI6w.
@@
[…]
@@
${}^{1}$epartment of Cognitive Robotics, Delft University of Technology, The Netherlands
@@ 1 Introduction
    Constrained function approximators. To ensure stability guarantees, methods often constrain the structure of their function approximators. For instance, some approaches necessitate invertibility [1, 2, 3], while others require positive or negative definiteness [4, 5, 6]. However, these methods do not enable full exploitation of modern Deep Neural Network (DNN) architectures, as these constraints are not typically present in DNNs. This limitation hinders their broader application in more complex models, where integrating these constraints is challenging. Furthermore, inherently constraining function approximators can overly restrict the range of solutions to which they can converge, resulting in less flexible models than necessary. This leads to suboptimal IL capabilities. In our context, this problem is known as the stability versus accuracy dilemma [7].
@@
    Euclidean assumption. Learned motions must integrate with the geometry of the space used to represent a robot’s state. For example, the end-effector of a manipulator should always reach the intended target in both position and orientation space. However, orientations are often represented in non-Euclidean spaces, such as $SO(3)$ or $\mathcal{S}^{3}$. This is because Euclidean representations like Euler angles may not always provide a continuous description of motions that are inherently continuous [^1] These limitations arise from singularities in these models, also known as gimbal lock, as well as non-unique representations. [8]. Such continuity is often a requirement for motions modeled as dynamical systems. Furthermore, the generalization capabilities of function approximations are compromised by non-continuous representations. To enable proper generalization, states that are close in the real world should be represented as closely related in the function approximator’s input, i.e., the state space representation.
@@
In this work, we present a DNN framework capable of learning accurate stable motions in state spaces with arbitrary geometries, without constraining the DNN’s architecture. To accomplish this, we introduce a novel loss function that repurposes the triplet loss, commonly used in deep metric learning literature [13, 14]. We prove that this loss imposes conditions on the DNN’s latent space that enforce the learned dynamical system to have a globally asymptotically stable equilibrium at the goal. To account for the geometry of the state space, it is sufficient to consider its corresponding metric during the computation of the loss, the choice of which depends on the specific task at hand. We validate our method in various settings, including Euclidean, non-Euclidean, first-order, and second-order motions. Additionally, real-world experiments that control the 7-dimensional pose (x-y-z position and quaternion orientation) of a robotic manipulator’s end effector demonstrate the method’s practical applicability and potential.
The rest of this paper offers a thorough discussion of these developments. Following a review of the relevant literature, we delve into foundational concepts and problem formulation. We then present the details of our methodology, provide a proof regarding the stability of the learned motion, and discuss the integration of the triplet loss function within the context of non-Euclidean state representations. We validate our approach through several experiments and conclude by considering potential directions for future research based on our findings.
@@ 2 Related Works
Three categories of works are relevant to this paper. First, there are papers that focus on learning stable motions, assuming these motions occur in Euclidean state spaces. Second, others explore learning motions in non-Euclidean state spaces, but do not consider the stability of the learned motions. Finally, a third set of papers addresses both learning motions in non-Euclidean state spaces and their stability.
In every case, works use either time-varying/non-autonomous or time-invariant/autonomous dynamical systems for motion modeling. In time-varying systems, evolution explicitly depends on time (or a phase). Conversely, time-invariant systems do not directly depend on time; instead, they rely on their time-varying input (i.e., the state of the system). The property of a system being time-invariant or not dictates the strategies we can use to ensure its stability. Thus, making a distinction between these systems is important. Notably, both types of formulations have been shown to be complementary in the context of IL [15, 16].
This work focuses on time-invariant dynamical systems. Consequently, although we consider both methodologies for motion learning, we delve deeper into the literature on time-invariant systems for IL.
@@ 2.1 Stability in Euclidean State Spaces
In previous work [9], we addressed this issue by applying tools from the deep metric learning literature [14]. In this context, we introduced CONDOR, which uses a contrastive loss to enforce stability in learned motions through the optimization process of a DNN. This approach proved effective in learning stable, accurate, and scalable motions from human demonstrations. However, this method presents two limitations. First, it requires the design of a stable and well-understood system for computing the contrastive loss (referred to as $f^{\mathcal{L}}$ in Sec. 3.3), which can significantly impact learning performance. Therefore, it would be advantageous to have a loss that does not depend on $f^{\mathcal{L}}$. Second, like the other methods introduced in this subsection, CONDOR is restricted to working in Euclidean state spaces.
In this work, we introduce a novel deep metric learning loss that ensures stability while addressing the limitations of [9], specifically, the need for the function $f^{\mathcal{L}}$ and the inability to learn motions in non-Euclidean manifolds.
@@ 2.2 Stability in Non-Euclidean State Spaces
These diffeomorphism-based methods are of particular interest to our work, as our method is grounded in similar concepts for achieving stability. In both approaches, we transfer the stability properties of a simple system in the latent space of a DNN to the dynamical system that models motion in task space. However, unlike these other methods, our approach learns this property, while the other works constrain the DNN structure to ensure its satisfaction. Moreover, our method allows us to effortlessly incorporate the geometrical aspects of the robot’s state space into the learned model, as this is factored into the optimization process of the DNN.
@@ 3.1 Dynamical Systems for Reaching Tasks
In this work, we model motions as solutions of nonlinear time-invariant dynamical systems represented by
$\dot{x}=f(x),$ (1)
where $x\in\mathcal{T}$, with $\mathcal{T}\subset\mathbb{R}^{n}$, is the state, and ${f:\mathcal{T}\to\mathbb{R}^{n}}$ is a differentiable function. Since we are interested in solving reaching tasks, we aim to construct systems with a globally asymptotically stable equilibrium at $x_{\mathrm{g}}$. This implies that
$\lim_{t\rightarrow\infty}||x_{\mathrm{g}}-x||=0,\hskip 14.22636pt\forall x\in\mathcal{T}.$ (2)
@@ 3.2 Problem Formulation
We consider a robot learning a reaching motion in the space $\mathcal{T}$ towards the goal state $x_{\mathrm{g}}\in\mathcal{T}$, based on a set of demonstrations $\mathcal{D}$. The robot is expected to imitate the behavior shown in the demonstrations while always reaching $x_{\mathrm{g}}$, regardless of its initial state. The dataset $\mathcal{D}$ contains $N$ trajectories $\tau_{i}$ showing the evolution of a dynamical system’s state $x_{t}$ from an initial state $x_{0}$ with $T$ samples obtained at different times $t$.
We assume that these demonstrations are drawn from the distribution $p^{*}(\tau)$, following the optimal dynamical system $f^{*}$ (according to the demonstrator’s judgment). On the other hand, the robot’s motion follows the parametrized system $f^{\mathcal{T}}_{\theta}$, inducing the distribution $p_{\theta}(\tau)$, where $\theta$ is the parameter vector. The objective is to find $\theta^{*}$ that minimizes the difference, expressed as the (forward) Kullback-Leibler divergence, between the trajectory distributions, while ensuring that $x_{\mathrm{g}}$ is a globally asymptotically stable equilibrium:
@@
s.t. $\lim_{t\rightarrow\infty}d(x_{\mathrm{g}},x_{t})=0,\forall x_{t}\in\mathcal{T}$ (3b)
where $\Theta$ is the parameter space of a DNN and $d(\cdot,\cdot)$ is a distance function.
@@ 3.3 Stability Conditions
In [9], the stability conditions were formulated to enforce stability in $f^{\mathcal{T}}_{\theta}$. The objective is to ensure that $f^{\mathcal{T}}_{\theta}$ inherits the stability properties from a simple and stable system, referred to as $f^{\mathcal{L}}$. Thus, if $f^{\mathcal{L}}$ is stable, $f^{\mathcal{T}}_{\theta}$ will also be stable. To achieve this, we require $f^{\mathcal{L}}$ to reside within the output of one of the hidden layers of the DNN that parameterizes $f^{\mathcal{T}}_{\theta}$. As such, we express the dynamical system $f^{\mathcal{T}}_{\theta}$ as a composition of two functions, $\psi_{\theta}$ and $\phi_{\theta}$,
@@
However, $f^{\mathcal{T}}_{\theta}$ is a standard DNN with $L$ layers. $\psi_{\theta}$ denotes layers $1,...,l$, and $\phi_{\theta}$ layers $l+1,...,L$. We define the output of layer $l$ as the latent space $\mathcal{L}\subset\mathbb{R}^{m}$. Note that $\mathcal{L}\subset\mathbb{R}^{m}$ implies the dimensionality of the vectors used to represent $\mathcal{T}$ and $\mathcal{L}$ does not need to be the same. Moreover, for simplicity, although we use the same $\theta$ notation for both $\psi_{\theta}$ and $\phi_{\theta}$, each symbol actually refers to a different subset of parameters within $\theta$. These subsets together form the full parameter set in $f^{\mathcal{T}}_{\theta}$.
Additionally, we need to introduce a third dynamical system. This system denotes the evolution in $\mathcal{L}$ of the states visited by $f^{\mathcal{T}}_{\theta}$ when mapped using $\psi_{\theta}$, which yields the relationship
$\dot{y}_{t}=f_{\theta}^{\mathcal{T}\to\mathcal{L}}(x_{t})=\frac{\partial\psi_{\theta}(x_{t})}{\partial t},$ (5)
$\forall y_{t}\in\mathcal{L}$, where $y_{t}$ corresponds to the latent variables that evolve according to $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$. Fig. 2 provides an example of the introduced dynamical systems.
@@
In this work, we reformulate these conditions and propose a novel loss for optimizing them. This adaptation endows the learning framework with increased flexibility, enabling it to tackle a wider array of problems (e.g., non-Euclidean state spaces) and achieve improved performance.
Figure 2: Example of trajectories generated by simulating the systems $f^{\mathcal{T}}_{\theta}$, $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ and $f^{\mathcal{L}}$ for different time instants using the forward Euler method. Importantly, the stability conditions are not met in this case, because $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ differs from $f^{\mathcal{L}}$.
@@ 3.4 Deep Metric Learning: the Triplet Loss
Commonly employed in the Deep Metric Learning literature, the triplet loss [13] has been utilized for learning and structuring latent state representations [14]. Its function is to cluster similar observations together and differentiate dissimilar ones within the latent space of a DNN. In this work, however, the triplet loss is used differently. Although we continue to use this loss to impose a certain structure on the DNN’s latent space, its goal is to do so in a specific manner that enforces stability in $f^{\mathcal{T}}_{\theta}$, achieved by satisfying the stability conditions.
@@
where $m\in\mathbb{R}_{>0}$ is the margin, $a$ is the anchor sample, $p$ the positive sample, and $n$ the negative sample This loss enforces positive samples to be at least $m$ distance closer to the anchor than the negative samples. Importantly, note that this is enough for the loss to become zero; it does not require the anchor and the positive sample to have the same value.

### v3

# PUMA: Deep Metric Imitation Learning for Stable Motion Primitives

@@ Abstract
Abstract— Imitation Learning (IL) is a powerful technique for intuitive robotic programming. However, ensuring the reliability of learned behaviors remains a challenge. In the context of reaching motions, a robot should consistently reach its goal, regardless of its initial conditions. To meet this requirement, IL methods often employ specialized function approximators that guarantee this property by construction. Although effective, these approaches come with some limitations: 1) they are typically restricted in the range of motions they can model, resulting in suboptimal IL capabilities, and 2) they require explicit extensions to account for the geometry of motions that consider orientations. To address these challenges, we introduce a novel stability loss function, drawing inspiration from the triplet loss used in the deep metric learning literature. This loss does not constrain the DNN’s architecture and enables learning policies that yield accurate results. Furthermore, it is not restricted to a specific state space geometry; therefore, it can easily incorporate the geometry of the robot’s state space. We provide a proof of the stability properties induced by this loss and empirically validate our method in various settings. These settings include Euclidean and non-Euclidean state spaces, as well as first-order and second-order motions, both in simulation and with real robots. More details about the experimental results can be found in: https://youtu.be/ZWKLGntCI6w.
@@
Rodrigo Pérez-Dattari1, Cosimo Della Santina1, and Jens Kober1
@@
1Department of Cognitive Robotics, Delft University of Technology, The Netherlands
@@ 1 Introduction
    Constrained function approximators. To ensure stability guarantees, methods often constrain the structure of their function approximators. For instance, some approaches necessitate invertibility [1, 2, 3], while others require positive or negative definiteness [4, 5, 6]. However, these methods do not enable the full exploitation of modern Deep Neural Network (DNN) architectures, as these constraints are not typically present in DNNs. This limitation hinders their broader application in more complex models, where integrating these constraints is challenging. Furthermore, inherently constraining function approximators can overly restrict the range of solutions to which they can converge, resulting in less flexible models than necessary. This leads to suboptimal IL capabilities. In our context, this problem is known as the stability versus accuracy dilemma [7].
@@
    Euclidean assumption. Learned motions must integrate with the geometry of the space used to represent a robot’s state. For example, the end-effector of a manipulator should always reach the intended target in both position and orientation space. However, orientations are often represented in non-Euclidean spaces, such as $SO(3)$ or $\mathcal{S}^{3}$. This is because Euclidean representations like Euler angles may not always provide a continuous description of motions that are inherently continuous [^1] This issue stems from singularities and non-unique representations. [8]. Such continuity is often a requirement for motions modeled as dynamical systems. Furthermore, the generalization capabilities of function approximations are compromised by non-continuous representations. To enable proper generalization, states that are close in the real world should be represented as closely related in the function approximator’s input, i.e., the state space representation.
@@
In this work, we present a DNN framework capable of learning accurate, stable motions in state spaces with arbitrary geometries without constraining the DNN’s architecture. To accomplish this, we introduce a novel loss function that repurposes the triplet loss, commonly used in deep metric learning literature [13, 14]. We prove that this loss imposes conditions on the DNN’s latent space that enforce the learned dynamical system to have a globally asymptotically stable equilibrium at the motion’s goal state (see Fig. 1). To account for the geometry of the state space, it is sufficient to consider its corresponding metric during the computation of the loss, the choice of which depends on the specific task at hand. We validate our method in various settings, including Euclidean, non-Euclidean, first-order, and second-order motions. Additionally, real-world experiments controlling the 6-dimensional pose (x-y-z position and unit quaternion orientation) of a robot manipulator’s end effector demonstrate the method’s practical applicability and potential.
The rest of this paper offers a thorough discussion of our developments. Following a review of the relevant literature, we delve into foundational concepts and problem formulation. We then present the details of our methodology, provide a proof regarding the stability of the learned motion, and discuss the integration of the triplet loss function within the context of non-Euclidean state representations. We validate our approach through several experiments and conclude by considering potential directions for future research based on our findings.
@@ 2 Related Works
Three categories of works are relevant to this paper. First, there are papers that focus on learning stable motions, assuming these motions occur in Euclidean state spaces. Second, others explore learning motions in non-Euclidean state spaces, but do not consider the stability of the learned motions. Finally, a third set of papers addresses both learning motions in non-Euclidean state spaces and their stability. Importantly,
In every case, works use either time-varying (non-autonomous) or time-invariant (autonomous) dynamical systems for motion modeling. In time-varying systems, evolution explicitly depends on time (or a phase). Conversely, time-invariant systems do not directly depend on time; instead, they rely on their time-varying input (i.e., the state of the system). The property of a system being time-invariant or not dictates the strategies we can use to ensure its stability. Thus, making a distinction between these systems is important. Notably, both types of formulations have been shown to be complementary in the context of IL [15, 16].
This work focuses on time-invariant dynamical systems. Consequently, although we consider both methodologies for motion learning, we delve deeper into the literature on time-invariant systems. Furthermore, while we concentrate on methods that ensure asymptotic stability, we acknowledge that there are studies imposing other conditions, such as via-point conditioning [17, 18], which can be significantly relevant in some robotic contexts.
@@ 2.1 Stability in Euclidean State Spaces
In previous work [9], we addressed this issue by employing tools from the deep metric learning literature [14]. There, we introduced CONDOR, which uses a contrastive loss to enforce stability in learned motions through the optimization process of a DNN. This approach proved effective in learning stable, accurate, and scalable motions from human demonstrations. However, this method presents two important limitations. First, it requires the design of a stable and well-understood system for computing the contrastive loss (referred to as $f^{\mathcal{L}}$ in Sec. 3.3), which can significantly impact learning performance. Second, similar to the other methods described in this subsection, CONDOR is limited to operating within Euclidean state spaces.
In this work, we introduce a novel deep metric learning loss that ensures stability while addressing the limitations of [9], specifically, the need for the function $f^{\mathcal{L}}$ and the inability to learn motions in non-Euclidean manifolds. Moreover, this loss requires weaker conditions for imposing stability, leading to, for instance, more robust performance when learning second-order dynamical systems.
@@ 2.2 Stability in Non-Euclidean State Spaces
These diffeomorphism-based methods are of particular interest to our work, as our method is grounded in similar concepts for achieving stability. In both approaches, we transfer the stability properties of a simple system in the latent space of a DNN to the dynamical system that models motion in task space. However, unlike these methods, our approach learns this property, whereas the other works constrain the DNN structure to ensure its satisfaction. Moreover, our method is not constrained to diffeomorphic solutions for transferring the stability properties of the simple system to task space. Lastly, our method allows us to seamlessly incorporate the geometrical aspects of the robot’s state space into the learned model, as this integration is factored into the DNN’s optimization process.
@@ 3.1 Dynamical Systems for Reaching Tasks
In this work, we model motions as nonlinear time-invariant dynamical systems represented by
$\dot{x}_{t}=f(x_{t}),$ (1)
where $x_{t}\in\mathcal{T}$, with $\mathcal{T}\subseteq\mathbb{R}^{n}$ being the task space, is the robot’s state, and ${f:\mathcal{T}\to\mathbb{R}^{n}}$ is a differentiable function. Additionally, the subscript $t$ indicates the time instance to which the state corresponds. Since we are interested in solving reaching tasks, we aim to construct systems with a globally asymptotically stable equilibrium at a goal state $x_{\mathrm{g}}$. This implies that
$\lim_{t\rightarrow\infty}||x_{\mathrm{g}}-x_{t}||=0,\hskip 14.22636pt\forall x_{t}\in\mathcal{T}.$ (2)
@@ 3.2 Problem Formulation
We consider a robot learning a reaching motion in the space $\mathcal{T}$ towards the goal state $x_{\mathrm{g}}\in\mathcal{T}$. Based on a set of demonstrations $\mathcal{D}$, the robot is expected to imitate the behavior shown in the demonstrations while always reaching $x_{\mathrm{g}}$, regardless of its initial state. The dataset $\mathcal{D}$ contains $N$ trajectories $\tau$. These trajectories show the evolution of a dynamical system’s state $x_{t}$ when starting from some initial condition $x_{0}$.
We assume that the demonstrations are drawn from an optimal distribution over trajectories, denoted as $p^{*}(\tau)$, that adheres to the optimal dynamical system $f^{*}$. Here, “optimal” refers to the behavior that the demonstrator deems best.
The robot’s motion follows the parametrized system $f^{\mathcal{T}}_{\theta}$, inducing the distribution $p_{\theta}(\tau)$, where $\theta$ is a parameter vector. Then, the objective is to find the vector $\theta^{*}$ that minimizes the difference, expressed as the (forward) Kullback-Leibler divergence, between $p_{\theta}(\tau)$ and $p^{*}(\tau)$, while ensuring that $x_{\mathrm{g}}$ is a globally asymptotically stable equilibrium:
@@
s.t. $\lim_{t\rightarrow\infty}d(x_{\mathrm{g}},x_{t})=0,\forall x_{t}\in\mathcal{T}.$ (3b)
Here, $\Theta$ is the parameter space of a DNN, and $d(\cdot,\cdot)$ is a distance function.
@@ 3.3 Stability Conditions
In [9], the stability conditions are formulated to enforce stability in $f^{\mathcal{T}}_{\theta}$. This is achieved by ensuring that $f^{\mathcal{T}}_{\theta}$ inherits the stability properties of a simple and stable system, referred to as $f^{\mathcal{L}}$. Thus, if $f^{\mathcal{L}}$ is asymptotically stable, $f^{\mathcal{T}}_{\theta}$ will also be asymptotically stable. To accomplish this, the system $f^{\mathcal{L}}$ is designed to define the evolution of a state $y_{t}$ with an initial condition derived from mapping $x_{0}$ to the output of a hidden layer in the DNN that parameterizes $f^{\mathcal{T}}_{\theta}$. As a result, the dynamical system $f^{\mathcal{T}}_{\theta}$ is expressed as a composition of two functions, $\psi_{\theta}$ and $\phi_{\theta}$,
@@
Here, $f^{\mathcal{T}}_{\theta}$ is a standard DNN with $L$ layers. $\psi_{\theta}$ denotes layers $1,...,l$, and $\phi_{\theta}$ layers $l+1,...,L$. We define the output of layer $l$ as the latent space $\mathcal{L}\subset\mathbb{R}^{m}$, which is where $y_{t}$ resides, i.e., $y_{t}\in\mathcal{L}$. Note that $\mathcal{L}\subset\mathbb{R}^{m}$ implies the dimensionality of the vectors used to represent $\mathcal{T}$ and $\mathcal{L}$ does not need to be the same. Moreover, for simplicity, although we use the same $\theta$ notation for both $\psi_{\theta}$ and $\phi_{\theta}$, each symbol actually refers to a different subset of parameters within $\theta$. These subsets together form the full parameter set in $f^{\mathcal{T}}_{\theta}$.
Additionally, we introduce a third dynamical system. This system denotes the evolution in $\mathcal{L}$ of the states visited by $f^{\mathcal{T}}_{\theta}$ when mapped using $\psi_{\theta}$, which yields the relationship
$\dot{y}_{t}=f_{\theta}^{\mathcal{T}\to\mathcal{L}}(x_{t})=\frac{\partial\psi_{\theta}(x_{t})}{\partial t}.$ (5)
Fig. 2 provides an example of the introduced dynamical systems.
@@
These conditions imply that if the system $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ behaves like $f^{\mathcal{L}}$, and any point other than $x_{\mathrm{g}}$ is excluded from mapping to the latent goal $y_{\mathrm{g}}$, then $f_{\theta}^{\mathcal{T}}$ is stable. Note that the latent goal is defined by the mapping $\psi_{\theta}(x_{\mathrm{g}})=y_{\mathrm{g}}$.
In this work, we reformulate these conditions and propose a novel way to optimize them. This adaptation endows the learning framework with increased flexibility, enabling it to tackle a wider array of problems (e.g., non-Euclidean state spaces) and achieve improved performance.
Figure 2: Example of trajectories generated by simulating the systems $f^{\mathcal{T}}_{\theta}$, $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ and $f^{\mathcal{L}}$ for different time instants. The stability conditions are not met in this case, as $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$ differs from $f^{\mathcal{L}}$.
@@ 3.4 Deep Metric Learning: the Triplet Loss
Commonly employed in the Deep Metric Learning literature, the triplet loss [13] has been utilized for learning and structuring latent state representations [14]. Its function is to cluster similar observations together and differentiate dissimilar ones within the latent space of a DNN. In this work, however, the triplet loss is used differently. Although we continue to use this loss to impose a certain structure on the DNN’s latent space, its purpose is to enforce the stability conditions, thereby ensuring that $f^{\mathcal{T}}_{\theta}$ is stable.
@@
where $m\in\mathbb{R}_{>0}$ is the margin, $a$ is the anchor sample, $p$ the positive sample, and $n$ the negative sample This loss enforces positive samples to be at least $m$ distance closer to the anchor than the negative samples. Notably, this is enough for the loss to become zero, i.e., it does not require the anchor and positive samples to have the same value.